6.5 pts Question 7 What is the final volume in ml) so that a 10 ml solution of HC10, at pH2 changes the ph to 020 010 OSO
Answer #1

HCO4 is a strong acid (complete dissociation)

HClO4 +H2O \rightarrow H3O+ + ClO4-

Therefore at pH =2

pH = -log[H3O+ ]

2= -log[H3O+ ]

[H3O+ ]= 10-2 M

Hence

[HClO4] = 10-2 M

At pH =4

pH = -log[H3O+ ]

4= -log[H3O+ ]

[H3O+ ]= 10-4 M

Hence

[HClO4] = 10-4 M

For dilution

M1V1= M2V2

Where M1 and V1 are initial molarity and volume and M2 and V2 are molarity and volume after dilution

For given problem

10-2 M X 10 mL = 10-4 M X V2

V2 = 1000 mL